OpenStudy (anonymous):

A square measures eight meters on each side. If each side is shortened by 1/4 of its original length, the perimeter of the new square will be ____ meters and its area will be ____ square meters.

OpenStudy (campbell_st):

I think you are missing some key information... the 2nd sentence If each side is shortened by???? or its original length...

OpenStudy (mathmale):

Did you mean "shortened to 1/3 of its original length?" Please refer to the original problem and find the multiplier in question. ...shortened to ??? of its original length

OpenStudy (anonymous):

1/4 @mathmale @campbell_st sorry

OpenStudy (anonymous):

8*0.25=?

OpenStudy (campbell_st):

ok... so you need to calculate the original perimeter P = 4 x 8 Area = 8 x 8 now find 3/4 of 8.... if you shorten a length by 1/4 its the same as using 3/4 of the length so what is the new length?

OpenStudy (campbell_st):

when you get the new length you can find the new perimeter and new area

OpenStudy (anonymous):

7.5

OpenStudy (campbell_st):

not quite.... its new length = (3 x 8)/4 or new length = 24/4 what do you think..?